The sum of a numerator and denominator is 15. If the denominator is increased by 1, the fractions become 1/3. Find the fraction.
 

Grade:8

2 Answers

Yuvraj Singh
2 years ago
Solution
Vijay
2 years ago
let , numerator of the given fraction be x ,
similarly for denominator be y.
According to question ,
 
numerator + denominator = x + y = 15 ,  ------ ( equation 1 )
Now , denominator is increased by 1 = y+1 ,
 
so new fraction { ( x )  / (y + 1 )  } = 1 /3 .
 
now cross multiplication , we get ,
3x = y + 1  ,
Now from equation  1 ,, we get -
3x= 15-x+1 ;
x = 4 
now from equation 1 , we get y = 11 .
 
so required fraction be 4/11.